OpenStudy (anonymous):

simplify √-49 divided by (3+4i)-(2-5i)

OpenStudy (anonymous):

please answer

OpenStudy (anonymous):

alright so we know √-49 = √-1*√49 = i7 this is what the numerator simplifies too, the bottom simplifies to 1 - i so we get 7i/1-i

OpenStudy (anonymous):

sorry the bottom is 1+9i

OpenStudy (kropot72):

\[\frac{i7}{1+i9}\]

OpenStudy (anonymous):

that not one of the answer though

OpenStudy (anonymous):

whats the answer they may have multiplied the top and the bottom by something

OpenStudy (anonymous):

is it -7/(i-9) ?

OpenStudy (anonymous):

7+7i divided by 2?

OpenStudy (anonymous):

63+7i divided by 82

OpenStudy (anonymous):

-7 + 7i divided by 2

OpenStudy (anonymous):

-63 + 7i divided by 82

OpenStudy (anonymous):

those are the choices...

OpenStudy (kropot72):

\[\frac{63+i7}{82}\]

OpenStudy (anonymous):

thats it kropot thanks!!

OpenStudy (kropot72):

You're welome :)

OpenStudy (anonymous):

can you explain it? im so confused

OpenStudy (anonymous):

they multiplied the top and the bottom by (1 - 9i)

OpenStudy (anonymous):

so in this case we got 7i/1+9i and now if multiply the top and the bottom by (1 - 9i) we get the bottom to be 1+81 =82 and the top to be 7i(1-9i) = 7i+63 so therefore the final answer is 7i+63/82 :)

OpenStudy (kropot72):

Sure.\[\frac{\sqrt{-49}}{(3+4i)-(2-5i)}=\frac{7i}{3+4i-2+5i}\] Do you follow?